# Spectral Small-Incremental-Entangling: Breaking Quasi-Polynomial Complexity Barriers in Long-Range Interacting Systems

### v1

@@ X Conclusion and perspective
In this work, we have aimed to establish universal principles governing the entanglement spectrum. The key notions introduced and developed are the SE strength (Definition 1) and the spectral SIE (Theorem 1). They provide a refined and quantitatively optimal framework for dynamical entanglement structure, thereby extending the celebrated SIE theorem to address richer many-body phenomena. As an application, we have analyzed long-range interacting systems and demonstrated that, contrary to previous expectations, the quasi-polynomial complexity of their simulation can in fact be improved to polynomial complexity. These main achievements are summarized in Fig. 1.
@@
    Unification of spectral and standard SIEs: The spectral SIE developed in this work has the distinct advantage that it bounds the generation rate of Rényi entanglement for $\alpha<1$, a feature that the standard SIE does not possess. This, in turn, allows one to control computational complexity in a more refined way. However, as discussed in Sec. V.2, the specialization of the spectral SIE to $\alpha=1$ does not necessarily recover the standard SIE. Moreover, as already emphasized in Ref. [34], the case $\alpha=1$ exhibits special properties that set it apart from $\alpha<1$. It therefore remains a major challenge to develop a unified theorem that fully integrates the spectral and the standard SIE frameworks into a comprehensive theory of dynamical entanglement growth.
    Related to this issue, our results show that the bound is sharp only for $\alpha=1/2$, but the optimal bound for $\alpha>1/2$ remains open.
@@
    Limits of operator approximability: Proposition 9 shows that small SE strength does not guarantee an efficient approximation of the time-evolution operator, thereby revealing a complexity separation between the full unitary and time-evolved states. However, once we allow the Schmidt rank to depend on the Hilbert space dimension $\mathcal{D}_{AB}$, it remains unresolved whether the required scaling is polynomial in $\mathcal{D}_{AB}$ or only logarithmic in $\mathcal{D}_{AB}$. Clarifying this point is crucial for achieving a more quantitative understanding of complexity separation in many-body quantum dynamics.
@@
    Sufficient conditions for operator-level low-rank approximation (Conjecture 2): This problem is closely related to the problem discussed above. While small SE strength is known to be insufficient to guarantee low-rank approximations at the operator level, it is conjectured that small $\alpha$-SE strength $\bar{J}_{\alpha}$ (Def. 2) might suffice. Proving this conjecture and determining the exact functional dependence $g_{\alpha}(D)$ would close a major gap between state-level and operator-level approximability.
@@
    Precise conditions for generalized area laws: The generalized area law currently relies on the assumption of boundary-adiabatic paths (Assumption 13). Whether this assumption is indeed indispensable or can be relaxed remains unknown. Determining the minimal assumptions under which area laws hold is essential for both condensed matter physics and quantum information theory.
@@
    Efficient algorithms for ground states and Gibbs states: Current results provide existence proofs of MPS (or purified MPS) approximations to gapped ground states and Gibbs states in 1D long-range interacting systems (Theorems 5 and 7). Yet explicit and efficient algorithms to actually construct such approximations are lacking. Developing such algorithms would make rigorous complexity results directly applicable to numerical simulation. More importantly, it would represent a genuine complexity-theoretic breakthrough, implying that the computational complexity of classical simulation of 1D long-range interacting systems lies in the class P.
@@ Acknowledgments
All the authors acknowledge the Hakubi projects of RIKEN. T. K. is supported by JST PRESTO (Grant No. JPMJPR2116), Exploratory Research for Advanced Technology (Grant No. JPMJER2302), and JSPS Grants-in-Aid for Scientific Research (No. JP23H01099, JP24H00071), Japan. Y. K. was supported by the JSPS Grant-in-Aid for Scientific Research (No. JP24K06909), Japan. This research was conducted during the internship of H. M. and C. R. at RIKEN, under the supervision of T. K.

### v2

@@ X Conclusion and perspective
In this work, we aimed to establish universal principles governing the entanglement spectrum. The key notions introduced and developed are the SE strength (Definition 1) and the spectral SIE (Theorem 1). They provide a refined and quantitatively optimal framework for dynamical entanglement structure, thereby extending the celebrated SIE theorem to address richer many-body phenomena. As an application, we have analyzed long-range interacting systems and demonstrated that, contrary to previous expectations, the quasi-polynomial complexity of their simulation can in fact be improved to polynomial complexity. These main achievements are summarized in Fig. 1.
@@
    Unification of spectral and standard SIEs: The spectral SIE developed in this work has the distinct advantage that it provides bounds on the generation rate of Rényi entanglement for $\alpha<1$, a feature not available in the standard SIE. This refinement allows for a more precise control of computational complexity. However, an important limitation arises: the spectral SIE, when specialized to $\alpha=1$, does not straightforwardly recover the standard SIE. As emphasized in Ref. [34], the case $\alpha=1$ exhibits special properties that set it apart from the regime $\alpha<1$. Consequently, a major open challenge is to develop a unified theorem that fully integrates the spectral and standard SIE frameworks into a comprehensive theory of dynamical entanglement growth. At the present stage of our understanding, our results show that the bound is sharp only for $\alpha=1/2$, whereas the optimal bound for $\alpha>1/2$ remains unknown.
@@
    Limits of operator approximability: Proposition 9 shows that small SE strength does not guarantee an efficient approximation of the time-evolution operator, thereby revealing a complexity separation between the full unitary operator and time-evolved states. However, once we allow the Schmidt rank to depend on the Hilbert space dimension $\mathcal{D}_{AB}$, it remains unresolved whether the required scaling is polynomial in $\mathcal{D}_{AB}$ or only logarithmic in $\mathcal{D}_{AB}$. Clarifying this point is crucial for achieving a more quantitative understanding of complexity separation in many-body quantum dynamics.
@@
    Sufficient conditions for operator-level low-rank approximation (Conjecture 2): This problem is closely related to the problem discussed above. While small SE strength is known to be insufficient to guarantee low-rank approximations at the operator level, it is conjectured that small $\alpha$-SE strength $\bar{{\mathcal{J}}}_{\alpha}$ (Def. 2) might suffice. Proving this conjecture and determining the exact functional dependence $g_{\alpha}(D)$ would close a major gap between state-level and operator-level approximability.
@@
    Precise conditions for generalized area laws: The generalized area law currently relies on the assumption of boundary-adiabatic paths (Assumption 13). It remains unknown whether this assumption is indeed indispensable or can be relaxed. Determining the minimal assumptions under which area laws hold is essential for both condensed matter physics and quantum information theory.
@@
    Time-efficient algorithms for ground states and Gibbs states: Current results provide existence proofs of MPS (or purified MPS) approximations to gapped ground states and Gibbs states in 1D long-range interacting systems (Theorems 5 and 7). However, explicit and efficient algorithms to actually construct such approximations are lacking. Developing such algorithms would make rigorous complexity results directly applicable to numerical simulation. More importantly, it would represent a genuine complexity-theoretic breakthrough, implying that the computational complexity of classical simulation of 1D long-range interacting systems lies in the class P.
@@ Acknowledgments
All the authors acknowledge the Hakubi projects of RIKEN. T. K. is supported by JST PRESTO (Grant No. JPMJPR2116), Exploratory Research for Advanced Technology (Grant No. JPMJER2302), and JSPS Grants-in-Aid for Scientific Research (No. JP23H01099, JP24H00071), Japan. Y. K. was supported by the JSPS Grant-in-Aid for Scientific Research (No. JP24K06909), Japan. The research of C.R. is supported by the DFG through the grant TRR 352 – Project-ID 470903074. This research was conducted during the internship of H. M. and C. R. at RIKEN, under the supervision of T. K.